OpenStudy (anonymous):

What is the standard form equation of the line passing through (5, –3) with a slope of –3?

OpenStudy (across):

\[y=m(x-x_1)+y_1\]

OpenStudy (across):

\[m=-3,\]\[x_1=5,\]\[y_1=-3.\]Do you know how to plug in?

OpenStudy (anonymous):

3x – y = –12 ???

OpenStudy (across):

Plug the values above into the equation above:\[y=-3(x-5)-3,\]\[y=-3x+15-3,\]\[y=-3x+12.\]

OpenStudy (anonymous):

so 3x – y = 12 ?

OpenStudy (paxpolaris):

Standard form: \[3x+y=12\]

OpenStudy (anonymous):

i hate these -_-